While being held captive on planet RAK, you notice that a rock takes 1.6 s to fall 2.5 m when you drop it. You also notice that a Newton scale reads 152 N when you step on it.  

a) what is your mass as determined on planet RAK?  

b) what would your mass be back on Earth?
Answer #1

The acceleration on RAK is

a = \frac{2s}{t^2} = \frac{2.5}{1.6^2 } = 0.97ms^{-2}

(a) The mass on the planet is

m = \frac{W}{a} = \frac{152}{0.97} = 156kg

(b) The mass on Earth would be the same.